# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 4 Enumeration formulae for self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}$
Using Magma, we provide the number $\mathfrak{S}_{4}(3;\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ of self-orthogonal codes of length $3$ and type $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ over $\mathscr{R}_{4,1}=GR(4,1)[y]/\langle y^{3}+2,2y\rangle$ in Table 1, the number $\mathfrak{S}_{5}(3;\lambda_{1},\lambda_{2},\ldots,\lambda_{5})$ of self-orthogonal codes of length $3$ and type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{5}\}$ over $\mathscr{R}_{5,1}=GR(4,1)[y]/\langle y^{3}+2,2y^{2}\rangle$ in Table 2, the number $\mathfrak{S}_{6}(2;\lambda_{1},\lambda_{2},\ldots,\lambda_{6})$ of self-orthogonal codes of length $2$ and type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{6}\}$ over $\mathscr{R}_{6,2}=GR(4,2)[y]/\langle y^{3}+2,2y^{3}\rangle$ in Table 3, and the number $\mathfrak{S}_{4}(4;\lambda_{1},\lambda_{2},\lambda_{,}\lambda_{4})$ of self-orthogonal codes of length $4$ and type $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ over $\mathscr{R}_{4,1}=GR(4,1)[y]/\langle y^{3}+2,2y\rangle$ in Table 4. These computed values agree exactly with Theorem 4.1.
From the above theorem, one can deduce Theorem 5.1 of Yadav and Sharma [32] on taking $\mathfrak{s}=1$ in the case when $e=\kappa$ is odd. Now, in the following theorem, we provide an explicit enumeration formula for the number $\mathfrak{U}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e}),$ where $e\geq 3.$
@@
One can deduce Theorem 5.3 of Yadav and Sharma [32] from the above theorem on taking $\mathfrak{s}=1$ in the case when $e=\kappa$ is odd.

### v2

@@ 4 Enumeration formulae for self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}$
Using Magma [6], we provide the number $\mathfrak{S}_{4}(3;\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ of self-orthogonal codes of length $3$ and type $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ over $\mathscr{R}_{4,1}=GR(4,1)[y]/\langle y^{3}+2,2y\rangle$ in Table 1, the number $\mathfrak{S}_{5}(3;\lambda_{1},\lambda_{2},\ldots,\lambda_{5})$ of self-orthogonal codes of length $3$ and type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{5}\}$ over $\mathscr{R}_{5,1}=GR(4,1)[y]/\langle y^{3}+2,2y^{2}\rangle$ in Table 2, the number $\mathfrak{S}_{6}(2;\lambda_{1},\lambda_{2},\ldots,\lambda_{6})$ of self-orthogonal codes of length $2$ and type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{6}\}$ over $\mathscr{R}_{6,2}=GR(4,2)[y]/\langle y^{3}+2,2y^{3}\rangle$ in Table 3, and the number $\mathfrak{S}_{4}(4;\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ of self-orthogonal codes of length $4$ and type $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ over $\mathscr{R}_{4,1}=GR(4,1)[y]/\langle y^{3}+2,2y\rangle$ in Table 4. These computed values agree exactly with Theorem 4.1.
Remark 4.2.
When $\kappa$ is odd, Theorem 5.1 of Yadav and Sharma [35] can be obtained as a special case of Theorem 4.1 by setting $\mathfrak{s}=1$ and $e=\mathtt{t}=\kappa.$
Now, in the following theorem, we provide an explicit enumeration formula for the number $\mathfrak{U}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e}),$ where $e\geq 3.$
@@
Remark 4.3.
When $\kappa$ is odd, Theorem 5.3 of Yadav and Sharma [35] follows as a direct consequence of Theorem 4.2 upon setting $\mathfrak{s}=1$ and $e=\mathtt{t}=\kappa.$
With the help of the enumeration formulae for self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m},$ obtained in Theorems 4.1 and 4.2, respectively, and by applying the classification algorithm described in [19, Sec. 9.6 and 9.7] and carrying out computations in Magma [6], one can obtain complete lists of inequivalent self-orthogonal and self-dual codes of a given type and length over $\mathscr{R}_{e,m}$ [34, 35, 36]. In the following example, we will now illustrate this by classifying all self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1}=GR(4,1)[y]/\langle y^{3}+2,2y\rangle,$ up to monomial equivalence. We will also explicitly determine a generator matrix for a representative of each equivalence class of such self-orthogonal codes.
Example 4.2.
From Table 1, we note that there are precisely $24$ distinct self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1}=GR(4,1)[y]/\langle y^{3}+2,2y\rangle.$ We will now apply the classification algorithm [19, Sec. 9.6 and 9.7] and carry out computations in Magma [6] to obtain all inequivalent self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1}.$
For a linear code $\mathcal{C}$ of length $n$ over $\mathscr{R}_{e,m},$ let $\mathcal{M}ono(\mathcal{C})$ denote the set of all monomial transformations that map $\mathcal{C}$ onto a self-orthogonal code over $\mathscr{R}_{e,m},$ and let $\mathcal{A}ut(\mathcal{C})$ denote the group of all monomial transformations that map the code $\mathcal{C}$ onto itself.
To begin with, we consider the self-orthogonal code $\mathcal{D}_{1}$ of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1}$ with a generator matrix
$\vskip-2.84526pt\begin{bmatrix}u&0&u+u^{3}\\ 0&u^{2}&u^{3}\end{bmatrix}$.
Using Magma [6], we compute $|\mathcal{M}ono(\mathcal{D}_{1})|=3072$ and $|\mathcal{A}ut(\mathcal{D}_{1})|=256.$ By the orbit–stabilizer principle [16, Th. 7.4], there are precisely $\frac{|\mathcal{M}ono(\mathcal{D}_{1})|}{|\mathcal{A}ut(\mathcal{D}_{1})|}=12$ distinct self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1}$ that are monomially equivalent to $\mathcal{D}_{1}.$
Using Magma [6] again, we further observe that the code $\mathcal{D}_{2}$ with a generator matrix
$\begin{bmatrix}u&0&u+u^{2}+u^{3}\\ 0&u^{2}&0\end{bmatrix}\vskip-2.84526pt$
is a self-orthogonal code of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1},$ which is monomially inequivalent to $\mathcal{D}_{1}.$ With the help of Magma, we compute $|\mathcal{M}ono(\mathcal{D}_{2})|=3072$ and $|\mathcal{A}ut(\mathcal{D}_{2})|=256.$ By the orbit–stabilizer principle [16, Th. 7.4] again, there are precisely $\frac{|\mathcal{M}ono(\mathcal{D}_{2})|}{|\mathcal{A}ut(\mathcal{D}_{2})|}=12$ distinct self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1},$ which are monomially equivalent to $\mathcal{D}_{2}.$
Hence, the total number of self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1}$ that are monomially equivalent to either $\mathcal{D}_{1}$ or $\mathcal{D}_{2}$ is $12+12=24,$ which coincides with the total number of such codes obtained from Table 1. Consequently, there are precisely two monomially inequivalent self-orthogonal codes of type $\{0,1,1,0\}$ and length $3$ over $\mathscr{R}_{4,1},$ viz. $\mathcal{D}_{1}$ and $\mathcal{D}_{2},$ with generator matrices
[…]
respectively.
@@ 5 Conclusion and future work
6 Acknowledgements
The authors acknowledge the research support provided by the Department of Science and Technology, India, under Grant no. DST/INT/RUS/RSF/P-41/2021 with TPN 65025.